Daily Events on 2024/02/15 (Thu)

last updated at  2024/02/15 10:32:01 JST


<< prev day

@

next day >>


Station

Pass No.
20240215-

Contact time (UT)

Max. Elev.
 (deg)

Events

AOS

LOS

USC20

0200

08:59:03

09:11:44

45.4

OP/OG upload
XRT: TI-CMD
EIS: TI-CMD
DR playback (by OP)

PRT1

2000

10:20:53

10:33:53

27.3

OP/OG upload, XRT:TI-CMD, EIS:TI-CMD (backup)
AOCS: Upload pointing table
EIS: Table Upload

USC20

0300

10:37:33

10:47:28

12.3

DR playback (by OP)
XRT: Table Upload

USC20

0400

20:59:56

21:11:27

15.7

DR playback (by OP)

USC20

0500

22:36:29

22:49:22

35.4

DR playback (by OP)

Flare list(LM) : http://www.lmsal.com/solarsoft/last_events/


DR information

OP main: 2024/02/15 11:16:00 - 2024/02/17 11:21:00 (UT)

Remaining amount
(prediction)

2786 Mbits @ OP_START ( 2024/02/15 11:16:00 )


DR usage: The DR would become empty on Feb 16 at 4 UT if all telescopes took no images until that time.


SOT

XRT

EIS

DR allocation [%]

54

23

23

DR allocation [Mbits]

4001

1704

1704

DR TLMLMT [Mbits]

2119 @OP_START
-> 5949

2513

2513


daily note

* Major Flare Watch on AR 13576 *

SOT:
- HOP 412 synoptic latitudinal scans (2/15) postponed to the weekend due to MFW
- HOP 393 eq. transition (2/17) postponed to the weekend due to MFW
- AR obs. (2/15, 2/16)

EIS:
- AR + flare observations

XRT:
- HOP 448 (Synoptic)
- AR Observations


[Maintenance Note (Yamagishi), for CO/CP/RT Tohban]
Some software packages will be updated and security patch will be installed on machines
'eisco', 'sotcows1', 'sotcows2', 'aocswin', 'higashirinkan', and 'ope-gw'
at 15:00-17:30(JST) on 15th February. When that's over, these machines will be rebooted.
Please do not use these machines during updating, and report if you have any trouble.


CHIEF
TOHBAN

SUB
TOHBAN

CP

SOT CO

XRT CO

EIS CO

EVENING

Kotanagi

Kotanagi

Tiwari (RCO)->Cruz (RCO)

Yoshimura (RCO)

Brooks (RCO)

MORNING


operation files

pointing file : cmdpln/20240215/re-point_202402150200.txt